A point charge (Q1) has a magnitude of -7 uC. A second charge (Q2) has a magnitude of -5.5 uC and is located 143 cm from the first charge. Find the magnitude of the force between the two charges.


1
Expert's answer
2022-02-28T10:38:49-0500

Given:

"q_1=-7.0*10^{-6}\:\rm C"

"q_2=-5.5*10^{-6}\:\rm C"

"r=1.43\:\rm m"


The Coulomb's law says


"F=k\frac{|q_1||q_2|}{r^2}""F=9*10^9*\frac{7.0*10^{-6}*5.5*10^{-6}}{1.43^2}=0.17 \:\rm N"